Recorded here is the frequency distribution of the blood types of 100 persons who have volunteered to donate blood at a plasma centre. Blood type O A B AB Total Frequency observed 40 44 10 6 100 Test the null hypothesis that the probability of the blood type O, A, B, AB is in the ratios 4:4:1:1. Use  = 0.05. Q


1
Expert's answer
2021-09-14T00:23:49-0400

"Χ^2 = \sum \frac{(O-E)^2}{E} \\

= \frac{(40-40)^2}{40} + \frac{(44-40)^2}{40} + \frac{(10-10)^2}{10} + \frac{(6-10)^2}{10} \\

= 0 + 0.4 + 0 + 1.6 \\

= 2.0"

P-value =CHISQ.DIST.RT(2, 3) = 0.5724

P-value > 0.05 => Accept the claim.

There is enough evidence to conclude that the probability of the blood type O, A, B, AB is in the ratios 4:4:1:1 at 0.05 level of significance.